# A note on thermal effects in non-linear models for plasma-based acceleration

### v1

@@ Abstract
We investigate the impact of a non-negligible background temperature on relativistic plasma wakefields generated when a beam of charged particles passes through a neutral plasma at rest. Our analysis focuses on the blowout regime, where the plasma response is highly non-linear: plasma electrons are radially blown out and expelled away from the propagation axis of the beam particles, creating a region (bubble) of ions without electrons. Our study builds upon earlier investigations for non-linear models of plasma wakefields developed neglecting plasma temperature. In the presence of a non-zero background temperature, we characterize the bubble in terms of its transversal and longitudinal sizes as a function of the temperature. Model predictions and parametrizations are studied in combination with PIC simulations, and correctly reproduce the temperature induced contraction of both the longitudinal and transverse bubble sizes.
@@ 1 Introduction
In plasma acceleration, a relativistic particle bunch or a short, intense laser pulse perturbs a plasma at rest, generating electromagnetic fields of order $\lesssim\mathcal{O}(100)\;\rm{GV/m}$ [1, 2] that can be used to accelerate and focus a witness of trailing beam particles properly injected in the plasma wake driven by the bunch/laser [3, 4]. These two scenarios are commonly referred to as plasma wakefield acceleration (PWFA) and laser wakefield acceleration (LWFA). Theories have been developed to understand wakefield properties in the linear regime [5, 6, 7]; however, when the plasma response becomes highly non-linear, a blowout regime is reached, wherein plasma electrons are radially blown out and expelled away from the propagating axis of the bunch/laser, due to the interaction with the charged particle bunch or ponderomotive force of the laser field. This forms a plasma cavity - commonly referred to as a bubble - that is devoid of electrons; expelled electrons are then attracted back towards the axis forming a plasma sheath around the bubble (see Fig.˜1 for a comprehensive sketch). In earlier seminal papers by Lu & coworkers [8, 9], a theory has been developed to describe the structure of these wakefields, based on an analysis in co-moving variables ($\xi=z-\beta_{b}ct$, with $\beta_{b}c$ the speed of the propagating bunch/laser and $z$ the coordinate along its propagation axis) and the quasi-static assumption: for a prescribed structure of the electron sheath, a non-linear differential equation for the bubble radius $r_{b}(\xi)$ and the inner electromagnetic fields could be derived. In the last 20 years, the model has been extended and improved in different ways: including the interaction of the electron bunch with the nonlinear plasma wave [10], considering different forms of the electron sheath [11, 12], including transverse plasma inhomogeneities [13, 14], considering multi-sheath plasma electron profiles [15], including bubble excitation [16] and also reformulating the theory based on the energy conservation law [17]. All these studies, however, have been performed in the cold limit, i.e., by neglecting thermal energy of plasma electrons. This choice is well motivated by the fact that the initial electron thermal energy in the plasma is expected to be of the order of $k_{b}T_{i}\sim 10\;\rm{eV}$ [18, 2], which is orders of magnitude smaller than the electron rest energy $m_{e}c^{2}=0.511$ MeV, resulting in an initial dimensionless temperature $\mu_{i}=k_{b}T_{i}/(m_{e}c^{2})\approx 10^{-5}$. Nevertheless, a complete theoretical framework cannot omit these contributions. Historically, thermal effects have been studied as a possible regularization mechanism for singularity of the wakefield near the wavebreaking point [19, 20, 21, 22]; importantly, recent research has significantly boosted renewed interest in plasma acceleration processes incorporating non-negligible thermal effects: heat accumulation may become a critical factor to control in high-repetition-rate operation [23]; in strongly nonlinear wakes, it has been shown that ion heating on the order of tens to hundreds of keV can counteract on-axis ion pinching, thereby enhancing wake stability [24]; then, following wave breaking, the energy released into the plasma was shown to drive electrons heating up to a few keV, influencing ion motion and enabling soliton formation [25, 26]. For positron acceleration, near-hollow channels with warm electrons (from sub-$\rm{keV}$ to ${\cal O}(10)\;\rm{keV}$) help suppress de-focusing forces and sustain stable propagation [27, 28, 29, 30]; lastly, light based diagnostics due to heating may offer new strategies to probe wake evolution [31]. This also prompted the development of novel numerical models to account for thermal effects [32, 33, 34], and motivates this current study that aims to revisit the model originally proposed by Lu & coworkers [8, 9] to include thermal effects. We derive the relevant equations for the bubble radius $r_{b}(\xi)$ in the presence of a non-negligible thermal background, which converge to the original equations provided by Lu in the limit $\mu_{i}\rightarrow 0$. These equations are then solved with a suitable boundary condition to account for non-zero thermal spread in the plasma electrons and the results are compared with particle in cell (PIC) numerical simulations.
@@ 2 Theory for warm electron bubble
In this section, we show all the calculations that lead to the non-linear ordinary differential equation for the bubble radius $r_{b}(\xi)$. The derivation shown here closely follows [8, 9], departing from it only for the introduction of thermal effects. For this reason, we will briefly give an overview of the general derivation, and highlight instead the additional terms emerging due to the presence of a non-zero temperature in plasma electrons.
For the rest of this work, we will assume dimensionless variables using as reference quantities the electron’s mass $m_{e}$, its absolute charge $e$ and the speed of light $c$. Consequently, relativistic momenta and energies will be normalized w.r.t. $m_{e}c$ and $m_{e}c^{2}$. Furthermore, number densities will always be given divided by the initial uniform rest number density of plasma electrons $n_{0}$, equal (due to charge neutrality) to the constant plasma ion’s number density $n_{i}$ – ions are considered immobile in our treatment. In this way, time is made dimensionless through $\omega_{p}^{-1}$, with $\omega_{p}=\sqrt{\frac{n_{0}e^{2}}{m_{e}\epsilon_{0}}}$ the cold plasma frequency, and lengths are made dimensionless through $k_{p}^{-1}$, with $k_{p}=\omega_{p}/c$ the cold plasma wave-number. Furthermore, for the sake of simplicity, we will focus here on a PWFA application, where the plasma response is induced by a bi-Gaussian beam of electrons moving at the speed of light in the $\hat{z}$ direction (from now on we assume $\beta_{b}=-1$) and with longitudinal/radial widths $\sigma_{z}/\sigma_{r}$. In co-moving coordinates, its normalized number density $n_{b}$ reads as:
@@
with $\alpha$ the normalized peak density amplitude, and $\xi_{0}$ its longitudinal co-moving origin. The whole procedure can easily be adapted to LWFA cases [8] where a laser source replaces this electron beam and the perturbing mechanism is due to ponderomotive force, since all the information pertaining the nature of the plasma perturbation is encoded in one single term of the $r_{b}(\xi)$ equation, as it is pointed out at the end of Section˜2.2.
@@ 2.1 Single electron equation of motion
Figure 1: Representative sketch coming from PIC simulations of the first electron depletion bubble generated by a relativistic electron bunch moving at the speed of light, in two different configurations of initial background plasma temperature. The observable represented in the figure is the number density, normalized w.r.t. the initial unperturbed value $n_{0}$, that consists of two contributions. The first is from the bi-Gaussian driving bunch $n_{b}$, characterized by normalized rms-sizes $\sigma_{r}=0.1$ and $\sigma_{z}=\sqrt{2}$, normalized peak amplitude $\alpha=84$, and propagating from right to left at the speed of light. The second contribution comes from the background plasma density $n_{e}$, initially set at uniform density $n_{0}=10^{16}cm^{-3}$ and uniform temperature $k_{b}T_{i}=0\;\rm{keV}$ (top half of the figure) and $k_{b}T_{i}=1\;\rm{keV}$ (bottom half of the figure). Also shown in gold is the curve $r_{b}(\xi)$, which marks the interface between the blowout region – where all plasma electrons have been expelled – and the outer region. Quantities $\ell_{\bot}$ and $\ell_{\parallel}$ measure respectively the transversal and longitudinal extents of this interface.
@@
where $\rho=1-n_{e}+n_{b}$ and $\mathbf{J}=-n_{e}\boldsymbol{\beta}+n_{b}\hat{\mathbf{z}}$ are the electromagnetic field sources, with $n_{e}$ the plasma electron number density and $\boldsymbol{\beta}$ its velocity. As is customary, one then adopts the co-moving variables $(x,y,z,t)\rightarrow(x,y,\xi=z-ct,t)$ and assumes the validity of the quasi-static approximation [36, 37] (QSA, $\partial_{t}\ll\partial_{\xi}$). Under this assumptions, Eq.˜2 becomes:
@@
The focus is now moved to the motion of a single plasma electron under the action of the electromagnetic potentials previously highlighted. Evolution equations for the relativistic normalized momentum $\mathbf{p}=\gamma\boldsymbol{\beta}$ and energy $E=\gamma$ of said electron are given by [38]:
@@
We remark that, at the time-scales of interest, the plasma can be regarded as collisionless. Accordingly, the only force acting on the electron in Eq.˜4 is the Lorentz force. Pressure contributions associated with a finite thermal spread in the electron momenta do not appear at the single-particle level: they emerge only as collective effects when the dynamics is described at the hydrodynamic scale. The imposition of the QSA, together with the choice of an ultra-relativistic bunch, allows to extract from Eq.˜4 a conservation equation in the form [38]:
@@
Eq.˜7 was also derived in [22, 28] when considering warm plasmas, under considerations similar to the ones presented here. Electron’s momentum $p_{z}^{(0)}$ and energy $E^{(0)}$ can be statistically sampled from an equilibrium Maxwell-Boltzmann distribution  [^1] In the relativistic regime, the correct momentum distribution is the Maxwell-Jüttner distribution [39]. However, for a gas at rest and sufficiently low temperatures ($\mu_{i}\ll 1$), this distribution reduces to the classical Maxwell-Boltzmann form that we are using in this work. with variance $k_{b}T_{i}$. By combining Eq.˜7 with the mass-shell condition $E^{2}-p_{z}^{2}-\mathbf{p}_{\perp}^{2}=1$, one can obtain a formula for both the energy and the relativistic longitudinal momentum of the electron:
@@
$E$ $=-\frac{1+\mathbf{p}_{\perp}^{2}+(\psi-E^{(0)}-p_{z}^{(0)})^{2}}{2(\psi-E^{(0)}-p_{z}^{(0)})}\;,$ (9)
@@
Eq.˜10 can be used to write the leading formula that will translate into a differential equation for the bubble trajectory $r_{b}(\xi)$. Under QSA, and considering the co-moving position of the electron $\xi=\xi[t,z(t)]$, one has:
@@ 2.2 Electromagnetic fields & sources
Figure 2: Numerical solutions of Eq.˜29, obtained using a fixed sheath thickness parameterization $\Delta=0.1r_{b}+1.0$ as in [8, 9]. Panel (a) shows a representative solution with an initial background plasma temperature of $k_{b}T_{i}=1\;\rm{keV}$. In the panel is shown an ensemble of electron trajectories (gray) obtained from 500 stochastic realizations of the initial conditions detailed in Sections˜3 and 34, and in orange the resulting average $\langle r_{b}\rangle$. Panel (b) compares this average trajectory (dashed orange line) against a cold-plasma realization of the model. While the mean trajectory remains close to the cold prediction, the trajectory bundle widens significantly, suggesting that thermal fluctuations contribute notably to the system’s behavior. This effect is quantified in panel (c), where the proposed fluctuation measure $\Delta r_{b}(\xi)$, evaluated at $\xi=11.0$ (highlighted in panel (a) by the black segment), is plotted as a function of the initial plasma temperature. The plot is in logarithmic scale.
Figure 3: Analysis of the first derivative $r_{b}^{\prime}$. Panel (a) shows an ensemble of derivatives’ trajectories (gray) obtained from 500500 stochastic a realization of the initial conditions, for an initial plasma temperature of $k_{b}T_{i}=1\;\rm{keV}$. The orange curve denotes the ensemble average $\langle r_{b}^{\prime}\rangle$. Panel (b) compares this average trajectory (dashed orange line) with the cold-plasma prediction, showing little to no difference between the two cases. Panel (c) shows $\Delta r_{b}^{\prime}(\xi)$, defined analogously to Eq.˜36.
Eq.˜15 constitutes the backbone of the equation we are seeking; to express the equation in a closed form, we need to identify the contribution due to the electromagnetic fields $A_{r},A_{z},\psi$. In order to do so, one has to write explicitly Eq.˜3 under axial symmetry:
$\frac{1}{r}\partial_{r}(rA_{r})$ $=\partial_{\xi}\psi\;,$ (16)
$\frac{1}{r}\partial_{r}(rJ_{r})$ $=\partial_{\xi}S\;,$ (17)
$-S$ $=\frac{1}{r}\partial_{r}(r\partial_{r}\psi)\;,$ (18)
$-J_{r}$ $=\frac{1}{r}\partial_{r}(r\partial_{r}A_{r})-\frac{1}{r^{2}}A_{r}\;,$ (19)
$-J_{z}$ $=\frac{1}{r}\partial_{r}(r\partial_{r}A_{z})\;.$ (20)
It is evident that most of the dynamics is governed by $S$, $J_{r}$ and $J_{z}$, which cannot be determined by first principles and therefore must be approximated via a model. Here, we will be using the so-called single-sheath model developed in [8, 9], which uses a radial stepwise approximation of $S$ along the blowout bubble. In the hypothesis of a complete blowout, where all the plasma electrons are expelled by the presence of a driving bunch, the authors in [8, 9] identify three main areas of interest:
@@
Here, $r_{b}(\xi)$ is the blowout radius at longitudinal position $\xi$, while $\Delta(\xi)$ denotes the thickness of the electron sheath surrounding the bubble. Note that when the plasma temperature is non-zero ($k_{b}T_{i}\neq 0$) the distinction among the three zones becomes less pronounced (see, for example, Figs.˜1 and 4), and a more refined parameterization (like, for example [15]) should be adopted. Nonetheless, for this preliminary study, we maintain the division into these three zones. Subsequent adjustments will be necessary to refine this approach for thermal scenarios.
A clear depiction of the electromagnetic sources within the bubble $(r\leq r_{b})$ can be achieved by considering that in a full blowout scenario, all plasma electrons have been expelled from this region, leaving only bunch electrons:
$\rho^{(in)}$ $=1-n_{b}\;,$ (22)
$J_{r}^{(in)}$ $=0\;,$ (23)
$J_{z}^{(in)}$ $=n_{b}\;.$ (24)
One can then substitute $J_{r}^{(in)}$ and $S^{(in)}=-\rho^{(in)}-J_{z}^{(in)}=-1$ into Eqs.˜18, 19 and 20 imposing no field divergence in $r=0$, the value $\psi_{0}=\psi(r=0)$, and using Eq.˜20 to determine integration constants:

### v2

@@ Abstract
We investigate the impact of a non-negligible background temperature on relativistic plasma wakefields generated when a beam of charged particles passes through a neutral plasma at rest. We focus on the blowout regime, wherein the plasma response is highly non-linear: plasma electrons are radially blown out and expelled away from the propagation axis of the beam particles, creating a region (bubble) of ions without electrons. Our study builds upon earlier investigations for non-linear models of plasma wakefields developed in the limit of zero background temperature (Lu et al., Phys. Rev. Lett. 96, 165002 (2006)). In the presence of a non-zero background temperature, we characterize the model and focus on its predictions for the bubble size and the electromagnetic fields inside the bubble. Model predictions are studied in combination with PIC simulations.
@@ 1 Introduction
In plasma acceleration, a relativistic particle bunch or a short, intense laser pulse perturbs a plasma at rest, generating electromagnetic fields of order $\lesssim\mathcal{O}(100)\;\rm{GV/m}$ [3, 19] that can be used to accelerate and focus a witness of trailing beam particles properly injected in the plasma wake driven by the bunch/laser [51, 12]. These two scenarios are commonly referred to as plasma wakefield acceleration (PWFA) and laser wakefield acceleration (LWFA). Theories have been developed to understand wakefield properties in the linear regime [24, 40, 31]; however, when the plasma response becomes highly non-linear, a blowout regime is reached, wherein plasma electrons are radially blown out and expelled away from the propagating axis of the bunch/laser, due to the interaction with the charged particle bunch or ponderomotive force of the laser field. This forms a ion cavity - commonly referred to as a bubble - that is devoid of electrons; expelled electrons are then attracted back towards the axis forming a plasma sheath around the bubble (see Fig.˜1 for a comprehensive sketch). The described regime is intrinsically non-linear, and lies beyond the reach of complete analytical theories. To accurately predict the blowout structure, and all of the observables of interest to wakefield acceleration, one must therefore resort to large scale kinetic simulations [12, 28]. In this context, Lu & co-workers [32, 30] proposed a phenomenological model that has become a cornerstone of the plasma-acceleration community, thanks to its strong predictive power and its ability to yield reliable scaling laws. The model, given some input parameters pertaining to the width of the electron sheath, provides an exact second-order differential equation for the bubble radius $r_{b}(\xi)$ with $\xi=z-\beta_{b}ct$ the co-moving coordinate ($\beta_{b}c$ is the speed of the propagating bunch/laser and $z$ the coordinate along its propagation axis; again, see Fig.˜1 for a comprehensive sketch). Solving this equation delivers reliable predictions of the shape of the ion cavity and the associated accelerating fields, without resorting to computationally demanding simulations. Important works in the field rely on this model to support their theoretical analyses. For instance, in high-quality acceleration experiments [29], it is used to compute the optimal beam loading for the trailing bunch that enables a flattened accelerating gradient and thus uniform energy gains for both electron [54, 55] and positron acceleration [59]; when intense laser pulses drive the wake, a self-injected electron beam may form during the plasma evolution [34, 37, 53] In this context, the model predicts the matching conditions required for stable laser propagation by determining the bubble shape [33]; in high repetition rate applications, where ion motion influences the plasma recovery process [7], the model has also been employed to support experimental data.
@@
In the last 20 years, the model has been extended and improved in different ways: including the interaction of the electron bunch with the nonlinear plasma wave [55], considering different forms of the electron sheath [57, 18], including transverse plasma inhomogeneities [17, 52], considering multi-sheath plasma electron profiles [8], including bubble excitation [16] and also reformulating the theory based on the energy conservation law [15]. All these studies, however, have been performed in the cold limit, i.e., by neglecting thermal energy of plasma electrons. This choice is well motivated by the fact that the initial electron thermal energy in the plasma is expected to be of the order of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k_{B}}T_{i}\sim 10\;\rm{eV}$ [1, 19], which is orders of magnitude smaller than the electron rest energy $m_{e}c^{2}=0.511$ MeV, resulting in an initial dimensionless temperature $\mu_{i}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k_{B}}T_{i}/(m_{e}c^{2})\approx 10^{-5}$. Nevertheless, a complete theoretical framework cannot omit these contributions. Historically, thermal effects have been studied as a possible regularization mechanism for singularity of the wakefield near the wavebreaking point [23, 41, 42, 21]; importantly, recent research has significantly boosted renewed interest in plasma acceleration processes incorporating non-negligible thermal effects: heat accumulation may become a critical factor to control in high-repetition-rate operation [7]; in strongly nonlinear wakes, it has been shown that ion heating on the order of tens to hundreds of keV can counteract on-axis ion pinching, thereby enhancing wake stability [14]; then, following wave breaking, the energy released into the plasma was shown to drive electrons heating up to a few keV, influencing ion motion and enabling soliton formation [58, 25]. For positron acceleration, near-hollow channels with warm electrons (from sub-$\rm{keV}$ to ${\cal O}(10)\;\rm{keV}$) help suppress de-focusing forces and sustain stable propagation [45, 9, 10, 5]; lastly, light based diagnostics due to heating may offer new strategies to probe wake evolution [26]. This also prompted the development of novel numerical models to account for thermal effects [36, 46, 47, 48], and motivates this current study that aims to revisit the model originally proposed by Lu & coworkers [32, 30] to include thermal effects. We derive the relevant equations for the bubble radius $r_{b}(\xi)$ in the presence of a non-negligible thermal background, which converge to the original equations provided by Lu in the limit $\mu_{i}\rightarrow 0$. These equations are then solved with a suitable boundary condition to account for non-zero thermal spread in the plasma electrons and the results are compared with particle in cell (PIC) numerical simulations.
@@
Figure 1: PIC snapshot of the first electron depletion bubble generated by an ultra-relativistic electron bunch propagating from right to left. The colormap represents the total normalized density ($n_{b}$ driver density, $n_{e}$ background plasma density), for two initial background temperatures: $k_{B}T_{i}=0.00\;\rm{keV}$ (top half of the figure) and $k_{B}T_{i}=1.00\;\rm{keV}$ (bottom half of the figure). The dashed white curve denotes the bubble boundary $r_{b}(\xi)$, the white solid lines denote its transverse and longitudinal extents, respectively $\ell_{\bot}$ and $\ell_{\parallel}$. The black line denotes the electron sheath $\Delta$.
@@ 2 Theory for warm electron bubble
In this section, we show all the calculations that lead to the non-linear ordinary differential equation for the bubble radius $r_{b}(\xi)$. The derivation shown here closely follows previous calculations [32, 30], departing from it only for the introduction of thermal effects. We remark that the physical framework in which the model is formulated does not rely on a fluid description, nor a mesoscopic one, but relies instead on tracking single representative plasma electrons, treating the dynamics from a Lagrangian, single-particle perspective. The derivation proceeds as follows: first, the equation of motion for a representative electron describing the radius of the bubble is examined under quasi-static, cylindrical, axially symmetric, and complete blowout assumptions. Then, the peculiar geometry of the bubble in the blowout regime is used to give a phenomenological estimate of the electromagnetic sources inside Maxwell’s equations, which are solved to obtain the electromagnetic fields, to be then plugged as a Lorentz force into the single electron equation of motion. The resulting equation is the Blowout Radius Equation, an ODE for $r_{b}(\xi)$ determining the bubble shape. In the cold limit, this recovers the Lu equation; in the warm case, additional thermal effects act on the initial conditions of the electron and modify force balance. [21, 9]. For the rest of this work, we will assume dimensionless variables using as reference quantities the electron’s mass $m_{e}$, its absolute charge $e$ and the speed of light $c$. Consequently, relativistic momenta and energies will be normalized w.r.t. $m_{e}c$ and $m_{e}c^{2}$. Furthermore, number densities will always be given divided by the initial uniform rest number density of plasma electrons $n_{0}$, equal (due to charge neutrality) to the constant plasma ion’s number density $n_{i}$: ions are considered immobile in our treatment. Consequently, current densities will be given divided by $cen_{0}$. In this way, time is made dimensionless through $\omega_{p}^{-1}$, with $\omega_{p}=\sqrt{\frac{n_{0}e^{2}}{m_{e}\epsilon_{0}}}$ the cold plasma frequency, and lengths are made dimensionless through $k_{p}^{-1}$, with $k_{p}=\omega_{p}/c$ the cold plasma wave-number. Furthermore, for the sake of simplicity, we will focus here on a PWFA application, where the plasma response is induced by a bi-Gaussian beam of electrons moving at the speed of light in the $\hat{z}$ direction (from now on we assume $\beta_{b}=-1$) and with longitudinal/radial widths $\sigma_{z}/\sigma_{r}$. In co-moving coordinates, its normalized number density $n_{b}$ reads as:
@@
with $\alpha$ the normalized peak density amplitude, and $\xi_{0}$ its longitudinal co-moving origin. The whole procedure can easily be adapted to LWFA cases [32] where a laser source replaces this electron beam and the perturbing mechanism is due to ponderomotive force, since all the information pertaining the nature of the plasma perturbation is encoded in one single term of the $r_{b}(\xi)$ equation, as it is pointed out at the end of Section˜2.2. Throughout this work, we will be expressing temperature using alternatively its dimensionless form (when convenient in formulas) and its physical form (mainly in figures, to give relevant points of comparison to the reader).
@@ 2.1 Single electron equation of motion
Figure 2: Numerical solutions of Eq.˜20 obtained using $\Delta=0.1r_{b}+1.0$. Panel a): representative warm plasma solution for $k_{B}T_{i}=1\;\rm{keV}$, showing 500500 stochastic realizations of the boundary-electron trajectories (gray) and their mean $\langle r_{b}\rangle$ (orange). Panel b): comparison between the mean warm trajectory (orange, dashed) and the cold-plasma solution. Panel c): measure of the fluctuations $\Delta r_{b}(\xi)$, evaluated at $\xi=11.0$ (black segment in panel a), plotted versus different background temperatures on a logarithmic scale.
@@
where $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho=1-n_{e}-n_{b}$} and $\mathbf{J}=-n_{e}\boldsymbol{\beta}+n_{b}\hat{\mathbf{z}}$ are the electromagnetic field sources, with $n_{e}$ the plasma electron number density and $\boldsymbol{\beta}$ its velocity. As is customary, one then adopts the co-moving variables $(x,y,z,t)\rightarrow(x,y,\xi=z-ct,t)$ and assumes the validity of the quasi-static approximation [49, 56] (QSA, $\partial_{t}\ll\partial_{\xi}$). Under this assumptions, Eq.˜2 becomes:
@@
The focus is now moved to the motion of a single plasma electron under the action of the electromagnetic potentials previously highlighted. Evolution equations for the relativistic normalized momentum $\mathbf{p}=\gamma\boldsymbol{\beta}$ and energy $E=\gamma$ (with $\gamma=(1-\boldsymbol{\beta}^{2})^{-1/2}$ the Lorentz factor) of said electron are given by [35]:
@@
We remark that, at the time-scales of interest, the plasma can be regarded as collisionless. Accordingly, the only force acting on the electron in Eq.˜4 is the Lorentz force. Pressure contributions associated with a finite thermal spread in the electron momenta do not appear at the single-particle level: they emerge only as collective effects when the dynamics is described at the hydrodynamic scale [47, 43, 44]. The imposition of the QSA, together with the choice of an ultra-relativistic bunch, allows to extract from Eq.˜4 a conservation equation in the form [35]:
@@
Eq.˜7 was also derived in [21, 9] when considering warm plasmas, under considerations similar to the ones presented here. Electron’s momentum $p_{z}^{(0)}$ and energy $E^{(0)}$ can be statistically sampled from an equilibrium Maxwell-Boltzmann distribution [^1] In the relativistic regime, the correct momentum distribution is the Maxwell-Jüttner distribution [22]. However, for a gas at rest and sufficiently low temperatures ($\mu_{i}\ll 1$), this distribution reduces to the classical Maxwell-Boltzmann form that we are using in this work. with variance ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k_{B}}T_{i}$. By combining Eq.˜7 with the mass-shell condition $E^{2}-p_{z}^{2}-\mathbf{p}_{\perp}^{2}=1$, one can obtain a formula for both the energy and the relativistic longitudinal momentum of the electron:
@@
$E$ $=-\frac{1+\mathbf{p}_{\perp}^{2}+(\psi-E^{(0)}-p_{z}^{(0)})^{2}}{2(\psi-E^{(0)}-p_{z}^{(0)})}\;,$
@@
Eq.˜9 can be used to write the leading formula that will translate into a differential equation for the bubble trajectory $r_{b}(\xi)$. Under QSA, and considering the co-moving position of the electron $\xi=\xi[t,z(t)]$, one has:
@@ 2.2 Electromagnetic fields & sources
Figure 3: Analysis of the first derivative $r_{b}^{\prime}$ for an initial plasma temperature of $k_{B}T_{i}=1.00\;\rm{keV}$. Panel a): ensemble of 500500 derivative trajectories (gray) and mean $\langle r_{b}^{\prime}\rangle$ (orange). Panel b): comparison between the mean warm trajectory (orange, dashed) and the cold-plasma solution. Panel c): fluctuations $\Delta r_{b}^{\prime}(\xi)$, defined analogously to Eq.˜27
Eq.˜14 constitutes the backbone of the equation we are seeking; to express the equation in a closed form, we need to identify the contribution due to the electromagnetic fields $A_{r},A_{z},\psi$. In order to do so, one has to solve explicitly Eq.˜3 under axial symmetry, after plugging in the corresponding electromagnetic sources $S$, $J_{r}$ and $J_{z}$. These cannot be determined by first principles and therefore must be approximated via a model. Here, we will be using the so-called single-sheath model developed in [32, 30], which uses a radial stepwise approximation of $S$ along the blowout bubble. In the hypothesis of a complete blowout, where all the plasma electrons are expelled by the presence of a driving bunch, the authors in [32, 30] identify three main areas of interest:
@@
Here, $r_{b}(\xi)$ is the blowout radius at longitudinal position $\xi$, while $\Delta(\xi)$ denotes the thickness of the electron sheath surrounding the bubble. Note that when the plasma temperature is non-zero (${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k_{B}}T_{i}\neq 0$) the distinction among the three zones becomes less pronounced (see, for example, Figs.˜1 and 4), and a more refined parameterization (like, for example [8]) should be adopted. Nonetheless, for this preliminary study, we maintain the division into these three zones. Subsequent adjustments will be necessary to refine this approach for thermal scenarios.
In a full blowout scenario, all plasma electrons have been expelled from within the bubble, leaving only bunch electrons, hence the electromagnetic sources are perfectly known therein; One can then obtain the electromagnetic potentials from Maxwell’s equations: